OpenStudy (anonymous):

MATH HELP PLEASE? :( Tania bought a glass pyramid from a gift store. The pyramid, with a height of 4 cm and a square base of side 6 cm, was packed in a box shaped as a cube of side 8 cm. The empty space in the box was packed with paper tissue. How many cubic centimeters of the box were stuffed with paper tissue? 368 cubic centimeters 464 cubic centimeters 192 cubic centimeters 480 cubic centimeters

OpenStudy (anonymous):

area of pyramid = half *area of base *height

OpenStudy (anonymous):

464? @THE_PROPHET

OpenStudy (anonymous):

@THE_PROPHET , area of pyramid????? are you sure?

OpenStudy (anonymous):

sorry its 1/3

OpenStudy (anonymous):

1/3*area of base *height

OpenStudy (anonymous):

it should be volume of pyramid

OpenStudy (anonymous):

look at your first comment @THE_PROPHET sir, u said are of pyramid=....... it should b volume i guess

OpenStudy (nikolas):

You need to find the volume of the box and subtract the volume of the pyramid. The volume of a pyramid is equal to bh*1/3 where b is the surface area of the pyramid's base (length times width, 6*6 in this case) and h is the height of the pyramid. The volume of a cube is equal to the product of its length, width, and height (8*8*8 here since all the sides are 8cm).

OpenStudy (anonymous):

i'M CONFUSED? Was The answer i picked right ?

OpenStudy (anonymous):

@THE_PROPHET @Nikolas @dg123

OpenStudy (nikolas):

Yes, the answer I got was 464. You find this by taking the volume of the box, 8*8*=512cm^3, and subtracting the volume of the pyramid, 6*6*4*(1/3)=48cm^3. 512-48= 464cm^3.

OpenStudy (anonymous):

answer would be 8^3-(4*6*6/3) =464,, hence the answer